OpenStudy (luckycoins888):

WILL FAN AND MEDALLL D: HALP MEH Mr. Fromm earned $100 by investing $4000 in a savings account earning simple interest for 6 months. Write an equation to find the annual interest rate. Solve. Show all work.

OpenStudy (luckycoins888):

@omgsomeonehelpmeplzD:

OpenStudy (luckycoins888):

Omg someone helllp

OpenStudy (luckycoins888):

OMG.

OpenStudy (luckycoins888):

WHERES ALL THE QUALIFIED HELPERs AT

OpenStudy (luckycoins888):

ffs

hartnn (hartnn):

so what do you know about simple interest? any formula?

OpenStudy (luckycoins888):

Not really..

hartnn (hartnn):

Simple interest = I = PNR where P = principal amount = 4000 I = 400 given N = number of years, here 6 months, so N = 1/2 R = interest rate is what we need to find :)

hartnn (hartnn):

so plug in the given values! only R is unknown

OpenStudy (luckycoins888):

OK, but i dont see how I would = 400

hartnn (hartnn):

`earned $100` its the interest that he earned by keeping $4000 in the bank. so, this $400 is actually the interest value = I

OpenStudy (luckycoins888):

Ohh ok

OpenStudy (luckycoins888):

So now what do we do e.e

hartnn (hartnn):

plug in known values! 400 = 4000 * (1/2) * R try to isolate R

OpenStudy (luckycoins888):

OK got it, now what?

hartnn (hartnn):

multiply 2 on both sides, and divide both sides by 4000 will give you, R = 400 * 2/ 4000 so what the value of R?

hartnn (hartnn):

are you there?

OpenStudy (luckycoins888):

Yes sorry i had to eat

hartnn (hartnn):

so what value of R did u get?

OpenStudy (luckycoins888):

i dont get the last step..

OpenStudy (luckycoins888):

is it 0.2?

hartnn (hartnn):

yes! good :) 0.2 or 20% both are same.

OpenStudy (luckycoins888):

Thanks!!!

hartnn (hartnn):

welcome ^_^